OpenStudy (anonymous):

Find the rate of change for each table,equation, or graph. How did you find the rate of change x 2 4 6 y 10 20 30

OpenStudy (anonymous):

please help asap

OpenStudy (anonymous):

alright so x=2, y=10 ... (2,10)... when x=4 y=20,(4,20) use these two in the slope formula \[\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}=\frac{20-10}{4-2}=\]